# uncompress

### v1

@@ 1.2 Scaling in language
The relationship between unique vocabulary and frequency is robust across many languages [piantadosi2014zipf_mechs_review, yu2018zipf_50langs_structural_pattern]; however, some work suggests exceptions to this robustness. Research using the Google N-gram corpus compared 7 languages’ more common lexicons and rarer lexicons, finding that in languages including Hebrew, Chinese, and Russian, rarer words are not Zipfian [petersen2012langs_cool_as_they_expand]. While finding good fit of lemmas in 4 languages, correlations between Zipf’s exponents for words and lemmas were not as robust for all languages, which could result from languages’ morphological complexity (English versus Finnish) [corral2015lemmas_zipf]. The type-token ratio divides vocabulary size (unique types) by total words to reflect text diversity; further work suggests that languages’ higher type-token ratio relates to higher morphological complexity and likely to produce many word forms [gutierrez_vasques2019morph_complexity]. Related, Zipfian and Heaps’ scaling do not hold for languages with limited dictionaries (e.g., Chinese, Japanese, and Korean) compared to English [lu2013zipf_lang_deviation]. These findings suggest that word segmentation and meaning encoding vary in languages and has scaling implications for languages producing many words. Thus, language and definition of a “word” (word-form) could drive steepness in scaling. [^1] Consider the morphological spectrum of languages from analytic to polysynthetic: polysynthetic languages generate novel word-forms continually (synthesizing concepts into one word), compared to analytic languages which tend to have fewer forms per word and relies on accompanying parts of speech for context [rijkhoff2002verbs].
@@ 3.1 Data sets
We use the CANDOR dataset [reece2023candor] and the Movie-Dialogs Corpus [danescu-niculescu-mizil2011ChameleonsImaginedConversations] for the present project. We reference two formulations of data from the Movie-Dialogs Corpus: “Movies (individual)” as meaning individual conversations among speakers whereas “Movies (grouped)” means all conversations in an entire movie have been joined to represent a script that only contains the movie’s speaking parts; we refer to these corpora as Movies (I) and Movies (G) for brevity. Movies (grouped) reflects a movie’s plot or story as conveyed via dialogue only. We provide dataset description and specific cleaning and subsetting methods in each corpus subsection. We distinguish between ‘word’ and ‘type’ when counts refer specifically to words versus inclusive of all tokenized ‘types’ such as punctuation.
@@ 3.1.1 CANDOR corpus
CANDOR data was collected in 2020 [reece2023candor]. The corpus contains $1,456$ unique participants. All participants were adults of ages 18 to 66, and some participants completed more than one conversation (see Appendix A1 for more details). Demographic data by participant included age, sex, race, education, and employment status. We remove one conversation because the raw data showed it was shortened because of technical difficulties. The final count of unique participants is $1,455$ ($1,431$ with demographic data). The conversation data contains $1,655$ English language conversations, $315,938$ conversation turns (utterances), and 9.29.2M word tokens.
@@ 3.1.2 Movie-Dialogs Corpus: Movies (I) and Movies (G)
The final corpus contains $2,397$ unique characters from $472$ unique movies throughout the release years 1927 to 2010. The conversation data includes $3,104$ conversations, $43,762$ utterances, and $485,389$ words. This corpus forms our Movies (individual) comparison group.
Movie conversations are very short compared to CANDOR’s naturalistic ones (see Table 2), so we formulate another exploration of Movie scaling to look at all conversations within one movie as a ‘conversation.’ Movie-Dialogs data does not ensure chronological ordering of all conversations within a movie, but prior work shows scaling replicates when compared to shuffled bodies of its own respective text [altmann2024statisticallaws]. The conjoined speaking parts formulation as another ‘conversation’ is conveying plot solely through dialogue from numerous participants. Still excluding conversations with null utterances or movies in non-fiction genres, this formulation allows us to include $8,481$ unique characters, $589$ movies (as $589$ conversations), totaling $288,125$ utterances and 33M words, and to have another comparison group called Movies (grouped).
@@ 3.2 Tools
All coding occurs in a Python 3.12 environment in Visual Studio Code, with libraries including NumPy, Pandas, and Matplotlib. See the project repository for dependencies and versions. Compute resources include an Apple M1 Macbook Pro with 32 GB of memory. Statistical analyses rely on SciPy for linear regression. Natural Language Processing (NLP) relies on regular expressions (re library) and NTLK for tokenization while Stanza’s pretrained English model [qi_stanza_2020] provides part-of-speech (POS) classification. Stanza’s model was trained on treebanks (containing linguistic dependency information), and Stanz’s POS processor utilizes sentence context in tag classification. Token labels come from universal part-of-speech tags [universaldependencies2021]. Stanza’s model shows high accuracy averages. [^4] Summarized from Stanza’s published model performance by language and treebank where the utilized treebanks are highly relevant to the current discourse types.
@@ 3.3.2 Statistical procedures
To explore the data, we use Pearson correlations, observe distributions, and analyze descriptive statistics. We also explore temporal measures of words and POS tags such as token interarrival times, burstiness, and memory [altmann2009beyond_word_freq, goh2008burstiness_memory]. Interarrival times are gaps between successive occurrences of types. Burstiness codifies the variation of interarrival times in a metric ($1<\text{B}<1$), with $\text{B}>0$ suggesting ‘bursty’ behavior (less regular) behavior while $\text{B}<0$ indicates more regular behavior. Memory is a correlation measure of interarrival times that codifies if these gaps are independent or show memory ($1<\text{M}<1$). $\text{M}>0$ indicates that similar length gaps follow each other (short and short or long and long). $\text{M}<0$ suggest more alternation between short and long gaps. Last, we calculate the type-token ratio (TTR) as a measure of lexical diversity or richness [gutierrez_vasques2019morph_complexity, rosillo_rodes2025entropy_ttr_gigaword] as vocabulary size (unique types) divided by total words where a value closer to 1 reflects high diversity.
@@ 4.1 Descriptive analyses
Figure 2: Words by utterances heatmap per corpus with Pearson correlation coefficients. The Movies (individual) corpus is subsetted to include conversations having a minimum of 10 utterances.
Corpora descriptive statistics ($\mu,~\sigma$) for number of words and utterances at the conversation level can be found in Table 2. CANDOR and Movies (G) result in the most similar structure for number of words while Movies (I) is distinctly much lower for both categories. Given the wide-ranging variance, we look at coefficients of variation ($\sigma/\mu$) to quantify the spread further. Of the six categories (words and utterances per corpus), number of words for Movies (I) ($0.59$) and Movies (G) ($0.48$) show the highest coefficients of variation (compared to $\text{CANDOR}=0.28$). Related, we look at the type-token ratio (TTR) and average across conversations’ TTRs at the corpus level. The corpus TTR distributions are normally distributed, and we report TTR as $\mu~(\sigma)$. Each corpus’s ratio appears distinct, with CANDOR (0.17(0.02)0.17~(0.02)) having the lowest diversity, followed by Movies (G) (0.24(0.07)0.24~(0.07)) and Movies (I) (0.64(0.09)0.64~(0.09)) with the highest diversity.
@@ 4.2 Temporal analyses
We calculate burstiness (B) and memory (M) along with multiple other comparison groups as we have not seen many textual results in this dual-metric format, excepting Ref. [goh2008burstiness_memory]. In Figure A6, we include B and M for: corpora, POS, literary works as a comparison (Frankenstein, Pride & Prejudice, and Moby Dick), and shuffled versions of each body denoted with $\ast$. We find differences for these literary texts compared to the 2 books reported in Ref. [goh2008burstiness_memory] which has burstiness and memory of effectively 0. Our literary comparisons cluster around $\text{B}=-0.09$, $\text{M}=0.35$. CANDOR and Movies (G) are close in burstiness to this literary cluster but lower in memory. Movies (I) (and its POS) shows the least memory (close to 0) but the most regularity (negative burstiness). Parts of speech differ more starkly from CANDOR and Movies (G) averages. Their interjections and function words are close to $\text{B}=0$ and $\text{M}=0$. The rest of the POS cluster around $\text{M}=0$ but trail into the lower left quadrant of less memory and less burstiness, which suggests more alternation between interarrival times and more regular behavior than random. Notably, all the Movies (I) POS are in the lower left quadrant. See Appendix A2 for temporal figures.
@@ 4.3 Regression analyses
We observe sublinear scaling and Heaps’ law similarly to prior works, e.g., [chacoma_tags_2020, linders2023zipf_spoken_dialog]. CANDOR reliably has longer conversations, and scaling exponents for individual conversations are normally distributed (Figure A1). For Movies (individual), conversations are less consistent and short. Each of the three corpora show CANDOR: $\beta=0.65\pm 0.0001$, Movies (I): $\beta=0.64\pm 0.003$, and Movies (G): $\beta=0.63\pm 0.0003$ (see Figure 4 for Heaps’ and corresponding rank regressions).
@@ 5.1 Presence of Heaps’ law in conversation
Our work identifies Heaps’ law scaling in naturalistic and fictional conversations. We find that scaling exponents are normally distributed for naturalistic conversations, consistent with multi-participant internet forums [kubo_2011]) and with past scaling behavior for spoken settings [abe_2021, bian2016scaling_like_lin]. The normal distribution of scaling exponents intriguingly offers a statistical artifact for future conversation research to replicate or to contrast to other forms of conversation or text. When comparing the three corpora, naturalistic and conjoined movie script dialogue are the most statistically similar across total words and type-token ratio, with both showing much lower variance around Heaps’ law exponents.
@@ 5.2 Part of speech differences by medium
Notably, our investigation of burstiness and memory yields a similar result. Movies (I) consistently show the most regularity as to when all of this dataset’s parts of speech occur, a regularity associated with function words [altmann2009beyond_word_freq]. This finding bolsters the thorough construction explanation of movie conversations. The idea of greater content clarity through function words also follows from movie dialogue possessing a stable and repetitive lexico-grammatical profile [zago2020film_discourse]. The short nature of Movies (I) conversations and high function word usage suggests that function words help structure meaning for viewers.
@@
We find that interjections’ scaling exponents are intermediate to those of open (noun, verb) and closed class (function words) scaling for the larger corpora (CANDOR and Movies (G)), which is consistent with interjections being a semi-open class [ameka1992interjections]. Movies (G) shows the highest scaling in interjections, but CANDOR shows the highest total and unique proportions of interjection use. This discrepancy between high scaling (type novelty) and usage (both new and repeating types recurring) could be a factor of document length, dual participants, and the nature of spontaneous conversation. Interjections serve an important feedback and reinforcement role in conversation between interlocutors [cooney2024naturalturn, boudin2024multimodal_feedback_in_convo] which is not functionally needed in prepared dialogue because it is more cohesively organized [martin2009teaching_convo_film, zago2020film_discourse], lending to lower use in movies [martin2010pos_bnc_movies]. It is apparent, at least, that the corpora rely on different interjections (Table A3), suggesting that different types of interjections are better suited to the purposes of conversation in each corpus; for instance, naturalistic speech shows more space-filling words (‘like’, ‘uh’, ‘um’) while we see more tags in fictional speech suggesting the introduction of concepts or contradiction (‘well’) and general use of ‘yes’ and ‘no’. Among the corpora, interjection usage is uniquely high in CANDOR, suggesting naturalistic conversation may rely on interjections as a connective supplement between participants and the meaning they attempt to convey (more-so than in movies [martin2010pos_bnc_movies]). In movies, actors are not trying to relate to each other in the same way as in spontaneous conversation (where interjections would be relational). Writing and physical movement are pre-scripted, so actors may not intuit or think to add vocal gestures despite attempting to act how humans act and sound. For the low and high interjection usage groups, we find no difference in group scaling within corpora. Interjections are an important part of building rapport, but they do not seem to have an additive effect on content in conversations. This null result suggests it may be more pro-social than informational, consistent with the conception of interjections as vocal gestures rather than speech [ameka1992interjections]. Another explanation is that linear regression is a limited method for delving into interjections. First, interjections are a difficult part of speech to parse automatically as its word-forms and recognizability highly varies. A POS tagger could miss instances, contributing to low counts and high uncertainty in a regression estimate. Second, interjections’ pro-social purpose provides more feedback to the listener than is contained in a single word expression, and analyzing unique and total words in a regression simply misses this context in many NLP analyses.
@@ 5.3 Conveying meaning
One reason for this language inefficiency could be the tendency to provide the most relevant information according to the delivery method—whether in-person or a writer imagining an audience’s reaction; under the maxim of quantity [grice1989wayofwords], the speaker has to judge how to provide maximally useful content so the listener comprehends the intent. For speakers to deem content maximally useful, speakers must adapt the dialogue to the circumstances in which they deliver it, considering the person receiving it, what layers of activity they can control (for example, speech, gestures, or writing direct messages), time of day, and other events taking place—circumstances broadly covering the dimensions of time and space [rijkhoff2002verbs]. Adapting to these circumstances also fits with Marconi’s two-strategy lexical competence explanation wherein inferential competence relates to structure and referential competence relates words to real entities [marconi1997LexicalCompetence] or representations thereof [varnier2000DiegoMarconiLexical]. The reliance of each corpus on particular strategies lets us consider how each corpus’s mean-making mechanisms may differ compared to standard speech. The different modalities’ inefficient parts of speech present a sort of fingerprint to make the most understanding in what is being conveyed, leaning more heavily on either lexical or referential competence; naturalistic and entire movie scripts’ reliance on content words (nouns and verbs) provides more pointers to the world and concepts. Single movie conversations’ reliance on verbs also makes use of this referential strategy but leans more heavily on lexical competence than the other corpora through function words. Speakers deliver naturalistic conversation using spoken utterances without a pre-determined structure to turns, allowing them the benefit of listeners’ interjections and other backchannels that helps speakers determine whether the listener understands the message. Movies in their entirety have the benefit of time and multimodality to convey their message. Instances of movie conversations have the challenge of ferrying audience from diegetic point to point, requiring that they maximize the space (perhaps the conversational floor) via more structure and specific verbiage. Actors do not have the benefit of interjection feedback from listeners to know the message is being perceived well.
@@
Overall, speakers or writers try to infer others’ mental states to communicate with them effectively [astington2005theory_of_mind_language]. With different affordances for communication, mediums shape communicative behavior via how speakers connect and know they are connecting.
@@
Speakers can leverage Theory of Mind and can maximize the space the conversational medium offers. Conversation itself is a communication tool in which we semantically encode and present our mental states, sometimes to convey complex things to other people [astington2005theory_of_mind_language]. Although much signal from prosody flattens in transcribed text, our work’s interjection results in transcribed spoken conversation suggests some mutual information with what is happening during spontaneous conversation [wolf2023prosody_text_redundancy]. In spontaneous conversation, speakers control the layers of activity of gestures and expression to convey their intended meaning. Interjections are both vocal gestures and, other times, expressions that enable listeners to show they have both perceived and understood the message—listeners are using theory of mind to demonstrate their understanding of the speaker. Authors of movie scripts access these layers but can also control numerous other elements to support communicative intent to help ‘listeners‘ understand the mind of the other. These elements could be lighting, sound, or costume to show emotions or state of mind. Writers in particular use mental models of people, situations, concepts, and etc. to represent concepts within speech. However, these mental models are communicated one way (without feedback and amendment to the message).

### v2

@@ 1.2 Scaling in language
The relationship between unique vocabulary and frequency is robust across many languages [piantadosi2014zipf_mechs_review, yu2018zipf_50langs_structural_pattern]; however, some work suggests exceptions to this robustness. Research using the Google N-gram corpus compared 7 languages’ more common lexicons and rarer lexicons, finding that in languages including Hebrew, Chinese, and Russian, rarer words are not Zipfian [petersen2012langs_cool_as_they_expand]. While finding good fit of lemmas in 4 languages, correlations between Zipf’s exponents for words and lemmas were not as robust for all languages, which could result from languages’ morphological complexity (English versus Finnish) [corral2015lemmas_zipf]. The type-token ratio divides vocabulary size (unique types) by total words to reflect text diversity; further work suggests that languages’ higher type-token ratio relates to higher morphological complexity and likely to produce many word forms [gutierrez_vasques2019morph_complexity]. Related, Zipfian and Heaps’ scaling do not hold for languages with limited dictionaries (e.g., Chinese, Japanese, and Korean) compared to English [lu2013zipf_lang_deviation]. These findings suggest that word segmentation and meaning encoding vary in languages and has scaling implications for languages producing many words. Thus, language and definition of a “word” (word-form) could drive steepness in scaling. [^1] Consider the morphological spectrum of languages from analytic to polysynthetic: polysynthetic languages generate novel word-forms continually (synthesizing concepts into one word), compared to analytic languages which tend to have fewer forms per word and rely on accompanying parts of speech for context [rijkhoff2002verbs].
@@ 3.1 Data sets
We use the Conversation: A Naturalistic Dataset of Online Recordings dataset (CANDOR) [reece2023candor] and the Movie-Dialogs Corpus [danescu-niculescu-mizil2011ChameleonsImaginedConversations] for the present project. We reference two formulations of data from the Movie-Dialogs Corpus: “Movies (individual)” as meaning individual conversations among speakers whereas “Movies (grouped)” means all conversations in an entire movie have been joined to represent a script that only contains the movie’s speaking parts; we refer to these corpora as Movies (I) and Movies (G) for brevity. Movies (grouped) reflects a movie’s plot or story as conveyed via dialogue only. We provide dataset description and specific cleaning and subsetting methods in each corpus subsection. We distinguish between ‘word’ and ‘type’ when counts refer specifically to words versus inclusive of all tokenized ‘types’ such as punctuation.
@@ 3.1.1 CANDOR corpus
CANDOR data were collected in 2020 [reece2023candor]. The corpus contains $1,456$ unique participants. All participants were adults of ages 18 to 66, and some participants completed more than one conversation (see Appendix A1 for more details). Demographic data by participant included age, sex, race, education, and employment status. We remove one conversation because the raw data showed it was shortened because of technical difficulties. The final count of unique participants is $1,455$ ($1,431$ with demographic data). The conversation data contains $1,655$ English language conversations, $315,938$ conversation turns (utterances), and 9.29.2M word tokens.
@@ 3.1.2 Movie-Dialogs Corpus: Movies (I) and Movies (G)
The final corpus contains $2,397$ unique characters from $472$ unique movies throughout the release years 1927 to 2010. The conversation data include $3,104$ conversations, $43,762$ utterances, and $485,389$ words. This corpus forms our Movies (individual) comparison group.
Movie conversations are very short compared to CANDOR’s naturalistic ones (see Table 2), so we formulate another exploration of Movie scaling to look at all conversations within one movie as a ‘conversation.’ Movie-Dialogs data do not ensure chronological ordering of all conversations within a movie, but prior work shows scaling replicates when compared to shuffled bodies of its own respective text [altmann2024statisticallaws]. The conjoined speaking parts formulation as another ‘conversation’ is conveying plot solely through dialogue from numerous participants. Still excluding conversations with null utterances or movies in non-fiction genres, this formulation allows us to include $8,481$ unique characters, $589$ movies (as $589$ conversations), totaling $288,125$ utterances and 33M words, and to have another comparison group called Movies (grouped).
@@ 3.2 Tools
All coding occurs in a Python 3.12 environment in Visual Studio Code, with libraries including NumPy, Pandas, and Matplotlib. See the project repository for dependencies and versions. Compute resources include an Apple M1 Macbook Pro with 32 GB of memory. Statistical analyses rely on SciPy for linear regression. NLP relies on regular expressions (re library) and NLTK for tokenization while Stanza’s pretrained English model [qi_stanza_2020] provides part of speech (POS) classification. Stanza’s model was trained on treebanks (containing linguistic dependency information), and Stanza’s POS processor utilizes sentence context in tag classification. Token labels come from universal part of speech tags [universaldependencies2021]. Stanza’s model shows high accuracy averages. [^4] Summarized from Stanza’s published model performance by language and treebank where the utilized treebanks are highly relevant to the current discourse types.
@@ 3.3.2 Statistical procedures
To explore the data, we use Pearson correlations, observe distributions, and analyze descriptive statistics. We also explore temporal measures of words and POS tags such as token interarrival times, burstiness, and memory [altmann2009beyond_word_freq, goh2008burstiness_memory]. Interarrival times are gaps between successive occurrences of types. Burstiness codifies the variation of interarrival times in a metric ($1<\text{B}<1$), with $\text{B}>0$ suggesting ‘bursty’ behavior (less regular) behavior while $\text{B}<0$ indicates more regular behavior. Memory is a correlation measure of interarrival times that codifies if these gaps are independent or show memory ($1<\text{M}<1$). $\text{M}>0$ indicates that similar length gaps follow each other (short and short or long and long). $\text{M}<0$ suggests more alternation between short and long gaps. Last, we calculate the type-token ratio (TTR) as a measure of lexical diversity or richness [gutierrez_vasques2019morph_complexity, rosillo_rodes2025entropy_ttr_gigaword] as vocabulary size (unique types) divided by total words where a value closer to 1 reflects high diversity.
@@ 4.1 Descriptive analyses
Figure 2: Words by utterances heatmap per corpus with Pearson correlation coefficients. The Movies (individual) corpus is subsetted to include conversations having a minimum of 10 utterances. Cell count coloring is in log10 space, with the global cell max =311=311.
Corpora descriptive statistics ($\mu,~\sigma$) for number of words and utterances at the conversation level can be found in Table 2. CANDOR and Movies (G) result in the most similar structure for number of words while Movies (I) is distinctly much lower for both categories. Given the wide-ranging variance, we look at coefficients of variation ($\sigma/\mu$) to quantify the spread further. Of the six categories (words and utterances per corpus), number of words for Movies (I) ($0.59$) and Movies (G) ($0.48$) shows the highest coefficients of variation (compared to $\text{CANDOR}=0.28$). Related, we look at the type-token ratio (TTR) and average across conversations’ TTRs at the corpus level. The corpus TTR distributions are normally distributed, and we report TTR as $\mu~(\sigma)$. Each corpus’s ratio appears distinct, with CANDOR (0.17(0.02)0.17~(0.02)) having the lowest diversity, followed by Movies (G) (0.24(0.07)0.24~(0.07)) and Movies (I) (0.64(0.09)0.64~(0.09)) with the highest diversity.
@@ 4.2 Temporal analyses
We calculate burstiness (B) and memory (M) along with multiple other comparison groups as we have not seen many textual results in this dual-metric format, excepting Ref. [goh2008burstiness_memory]. In Figure A6, we include B and M for: corpora, POS, literary works as a comparison (Frankenstein, Pride & Prejudice, and Moby Dick), and shuffled versions of each body denoted with $\ast$. We find differences for these literary texts compared to the 2 books reported in Ref. [goh2008burstiness_memory] which have burstiness and memory of effectively 0. Our literary comparisons cluster around $\text{B}=-0.09$, $\text{M}=0.35$. CANDOR and Movies (G) are close in burstiness to this literary cluster but lower in memory. Movies (I) (and its POS) shows the least memory (close to 0) but the most regularity (negative burstiness). Parts of speech differ more starkly from CANDOR and Movies (G) averages. Their interjections and function words are close to $\text{B}=0$ and $\text{M}=0$. The rest of the POS cluster around $\text{M}=0$ but trail into the lower left quadrant of less memory and less burstiness, which suggests more alternation between interarrival times and more regular behavior than random. Notably, all the Movies (I) POS are in the lower left quadrant. See Appendix A2 for temporal figures.
@@ 4.3 Regression analyses
We observe sublinear scaling and Heaps’ law similarly to prior works, e.g., [chacoma_tags_2020, linders2023zipf_spoken_dialog]. CANDOR reliably has longer conversations, and scaling exponents for individual conversations are normally distributed (Figure A1). For Movies (individual), conversations are less consistent and short. The three corpora show CANDOR: $\beta=0.65\pm 0.0001$, Movies (I): $\beta=0.64\pm 0.003$, and Movies (G): $\beta=0.63\pm 0.0003$ (see Figure 4 for Heaps’ and corresponding rank regressions).
@@ 5.1 Presence of Heaps’ law in conversation
Our work identifies Heaps’ law scaling in naturalistic and fictional conversations. We find that scaling exponents are normally distributed for naturalistic conversations, consistent with multi-participant internet forums [kubo_2011] and with past scaling behavior for spoken settings [abe_2021, bian2016scaling_like_lin]. The normal distribution of scaling exponents intriguingly offers a statistical artifact for future conversation research to replicate or to contrast to other forms of conversation or text. When comparing the three corpora, naturalistic and conjoined movie script dialogue are the most statistically similar across total words and type-token ratio, with both showing much lower variance around Heaps’ law exponents.
@@ 5.2 Part of speech differences by medium
Notably, our investigation of burstiness and memory yields a similar result. Movies (I) consistently shows the most regularity as to when all of this dataset’s parts of speech occur, a regularity associated with function words [altmann2009beyond_word_freq]. This finding bolsters the thorough construction explanation of movie conversations. The idea of greater content clarity through function words also follows from movie dialogue possessing a stable and repetitive lexico-grammatical profile [zago2020film_discourse]. The short nature of Movies (I) conversations and high function word usage suggests that function words help structure meaning for viewers.
@@
We find that interjections’ scaling exponents are intermediate to those of open (noun, verb) and closed class (function words) scaling for the larger corpora (CANDOR and Movies (G)), which is consistent with interjections being a semi-open class [ameka1992interjections]. Movies (G) shows the highest scaling in interjections, but CANDOR shows the highest total and unique proportions of interjection use. This discrepancy between high scaling (type novelty) and usage (both new and repeating types recurring) could be a factor of document length, dual participants, and the nature of spontaneous conversation. Interjections serve an important feedback and reinforcement role in conversation between interlocutors [cooney2024naturalturn, boudin2024multimodal_feedback_in_convo] which is not functionally needed in prepared dialogue because it is more cohesively organized [martin2009teaching_convo_film, zago2020film_discourse], lending to lower use in movies [martin2010pos_bnc_movies]. It is apparent, at least, that the corpora rely on different interjections (Table A3), suggesting that different types of interjections are better suited to the purposes of conversation in each corpus; for instance, naturalistic speech shows more space-filling words (‘like’, ‘uh’, ‘um’) while we see more tags in fictional speech suggesting the introduction of concepts or contradiction (‘well’) and general use of ‘yes’ and ‘no’. Among the corpora, interjection usage is uniquely high in CANDOR, suggesting naturalistic conversation may rely on interjections as a connective supplement between participants and the meaning they attempt to convey (more-so than in movies [martin2010pos_bnc_movies]). In movies, actors are not trying to relate to each other in the same way as in spontaneous conversation (where interjections would be relational). Writing and physical movement are pre-scripted, so actors may not intuit or think to add vocal gestures despite attempting to act how humans act and sound. For the low and high interjection usage groups, we find no difference in group scaling within corpora. Interjections are an important part of building rapport, but they do not seem to have an additive effect on content in conversations. This null result suggests it may be more pro-social than informational, consistent with the conception of interjections as vocal gestures rather than speech [ameka1992interjections]. Another explanation is that linear regression is a limited method for delving into interjections. First, interjections are a difficult part of speech to parse automatically as its word-forms and recognizability highly vary. A POS tagger could miss instances, contributing to low counts and high uncertainty in a regression estimate. Second, interjections’ pro-social purpose provides more feedback to the listener than is contained in a single word expression, and analyzing unique and total words in a regression simply misses this context in many NLP analyses.
@@ 5.3 Conveying meaning
One reason for this language inefficiency could be the tendency to provide the most relevant information according to the delivery method—whether in-person or a writer imagining an audience’s reaction; under the maxim of quantity [grice1989wayofwords], the speaker has to judge how to provide maximally useful content so the listener comprehends the intent. For speakers to deem content maximally useful, speakers must adapt the dialogue to the circumstances in which they deliver it, considering the person receiving it, what layers of activity they can control (for example, speech, gestures, or writing direct messages), time of day, and other events taking place—circumstances broadly covering the dimensions of time and space [rijkhoff2002verbs]. Adapting to these circumstances also fits with Marconi’s two-strategy lexical competence explanation wherein inferential competence relates to structure and referential competence relates words to real entities [marconi1997LexicalCompetence] or representations thereof [varnier2000DiegoMarconiLexical]. The reliance of each corpus on particular strategies lets us consider how each corpus’s mean-making mechanisms may differ compared to standard speech. The different modalities’ inefficient parts of speech present a sort of fingerprint to make the most understanding in what is being conveyed, leaning more heavily on either inferential or referential lexical competence; naturalistic and entire movie scripts’ reliance on content words (nouns and verbs) provides more pointers to the world and concepts (referential competence). Single movie conversations’ reliance on verbs also makes use of this lexical strategy but leans more heavily on inferential competence than the other corpora through function words. Speakers deliver naturalistic conversation using spoken utterances without a pre-determined structure to turns, allowing them the benefit of listeners’ interjections and other backchannels that help speakers determine whether the listener understands the message. Movies in their entirety have the benefit of time and multimodality to convey their message. Instances of movie conversations have the challenge of ferrying audience from diegetic point to point, requiring that they maximize the space (perhaps the conversational floor) via more structure and specific verbiage. Actors do not have the benefit of interjection feedback from listeners to know the message is being perceived well.
@@
Overall, speakers or writers try to infer others’ mental states to communicate with them effectively [astington2005theory_of_mind_language]. With different affordances for communication, mediums shape communicative behavior via how speakers connect and know they are connecting [mcluhan1964medium].
@@
Speakers can leverage Theory of Mind and can maximize the space the conversational medium offers. Conversation itself is a communication tool in which we semantically encode and present our mental states, sometimes to convey complex things to other people [astington2005theory_of_mind_language]. Although much signal from prosody flattens in transcribed text, our work’s interjection results in transcribed spoken conversation suggest some mutual information with what is happening during spontaneous conversation [wolf2023prosody_text_redundancy]. In spontaneous conversation, speakers control the layers of activity of gestures and expression to convey their intended meaning. Interjections are both vocal gestures and, other times, expressions that enable listeners to show they have both perceived and understood the message—listeners are using theory of mind to demonstrate their understanding of the speaker. Authors of movie scripts access these layers but can also control numerous other elements to support communicative intent to help ‘listeners‘ understand the mind of the other. These elements could be lighting, sound, or costume to show emotions or state of mind. Writers in particular use mental models of people, situations, concepts, and etc. to represent concepts within speech. However, these mental models are communicated one way (without feedback and amendment to the message).